Apply De Moivre’s formula to express cos 4θ and sin 4θ in terms of cos θ and sin θ.


1
Expert's answer
2021-12-27T03:23:46-0500

Let us apply De Moivre’s formula to express "\cos 4θ" and "\sin 4θ" in terms of "\cos θ" and "\sin θ." It follows from De Moivre’s formula that "(\cos\theta+i\sin\theta)^4=\cos4\theta+i\sin4\theta," where "i^2=-1."

On the other hand, the Binomial Theorem implies that "(\cos\theta+i\sin\theta)^4
\\=\cos^4\theta+4\cos^3\theta\cdot i\sin\theta+6\cos^2\theta\cdot i^2\sin^2\theta
+4\cos\theta\cdot i^3\sin^3\theta+i^4\sin^4\theta
\\=\cos^4\theta+4i\cos^3\theta\cdot \sin\theta-6\cos^2\theta\cdot \sin^2\theta
-4i\cos\theta\cdot \sin^3\theta+\sin^4\theta
\\=\cos^4\theta-6\cos^2\theta\cdot \sin^2\theta
+\sin^4\theta+i(4\cos^3\theta\cdot \sin\theta-4\cos\theta\cdot \sin^3\theta)."


Taking into account that two complex numbers are equal if and only if their real parts are equal and their imaginary parts are equal, we conclude that


"\cos4\theta=\cos^4\theta-6\cos^2\theta\cdot \sin^2\theta
+\sin^4\theta,
\\
\sin4\theta=4\cos^3\theta\cdot \sin\theta-4\cos\theta\cdot \sin^3\theta."